Licence / Master Science de la matière  Stage 2019-2020
École Normale Supérieure de Lyon  Antoine MARIE
Université Claude Bernard Lyon I  M1 Chimie

Pertubation theories in the complex plane

Abstract :

Keywords : perturbation theory, Hartree-Fock, spherium, exceptional points

Internship supervisor: Pierre-François LOOS

loos@irsamc.ups-tlse.fr/ tél. (+33) 5 61 55 73 39

Laboratoire de Chimie et Physique Quantiques 118, route de Narbonne 31062 Toulouse - France

https://www.irsamc.ups-tlse.fr/

July 10, 2020

Perturbation theory in the complex plane  Antoine MARIE

Acknowledgments Contents

1 Introduction  3
2 Perturbation theory  3
3 Historical overview  5
3.1 Behavior of the Møller-Plesset series . . . . . . . . . . . . . . . . . . . . . . . . . . . .  5
3.2 The singularity structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  7
3.3 The physics of quantum phase transition . . . . . . . . . . . . . . . . . . . . . . . . . .  8
4 The spherium model  8

Perturbation theory in the complex plane  Antoine MARIE

1 Introduction

It has always been of great importance for theoretical chemists to better understand excited states and their properties because processes involving excited states are ubiquitous in nature (physics, chemistry and biology). One of the major challenges is to accurately compute excited states energies. Plenty of methods have been developed to this aim and each of them have its own qualities but also its own flaws. The fact that none of all those methods is succesfull for every molecule in every geometry encourage chemists to continue the development of new methodologies to get accurate energies and to try to understand deeply why each methods fails or not with each molecule. All those methods rely on the notion of quantised energy levels of Hermitian quantum mechanics. In quantum chemistry, the ordering of the energy levels represents the different electronic states of a molecule, the lowest being the ground state while the higher ones are the so-called excited states. And we need methods to get accurately how those states are ordered. Within this quantised paradigm, electronic states look completely disconnected from one another. However, one can gain a different perspective on quantisation if one extends quantum chemistry into the complex domain. In a non-Hermitian complex picture, the energy levels are sheets of a more complicated topological manifold called Riemann surface, and they are smooth and continuous analytic continuation of one another. In other words, our view of the quantised nature of conventional Hermitian quantum mechanics arises only from our limited perception of the more complex and profound structure of its non-Hermitian variant. Therefore, by analytically continuing the energy E(λ) in the complex domain (where λ is a coupling parameter), one can smoothly connect the ground and excited states of a molecule. This connection is possible because, by extending real numbers to the complex domain, one loses the ordering property of real numbers. Hence, one can interchange electronic states away from the real axis, as the concept of ground and excited states has been lost. Amazingly, this smooth and continuous transition from one state to another has been recently realized experimentally in physical settings such as electronics, microwaves, mechanics, acoustics, atomic systems and optics. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] Exceptional points (EPs) [19, 20, 21, 22] are non-Hermitian analogs of conical intersections (CIs) [23] where two states become exactly degenerate. CIs are ubiquitous in non-adiabatic processes and play a key role in photo-chemical mechanisms. In the case of auto-ionizing resonances, EPs have a role in deactivation processes similar to CIs in the decay of bound excited states. Although Hermitian and non-Hermitian Hamiltonians are closely related, the behavior of their eigenvalues near degeneracies is starkly different. For example, by encircling non-Hermitian degeneracies at EPs leads to an interconversion of states, and two loops around the EP are necessary to recover the initial energy. Additionally, the wave function picks up a geometric phase (also known as Berry phase [24]) and four loops are required to recover the starting wave function. In contrast, encircling Hermitian degeneracies at CIs introduces only a geometric phase while leaving the states unchanged. More dramatically, whilst eigenvectors remain orthogonal at CIs, at non-Hermitian EPs the eigenvectors themselves become equivalent, resulting in a self-orthogonal state. [25] More importantly here, although EPs usually lie off the real axis, these singular points are intimately related to the convergence properties of perturbative methods and avoided crossing on the real axis are indicative of singularities in the complex plane. [26, 27]

2 Perturbation theory

Within perturbation theory, the Schrödinger equation is usually rewritten as

H(λ)Ψ = (H^{(0)} + λV)Ψ(λ) = E(λ)Ψ,  (1)

with

V = H − H^{(0)}  (2)

The energy can then be written as a power series of λ

E(λ) = ∑_{k=0}^{∞} λ^{k}E^{(k)}  (3)

where λ is a coupling parameter set equal to 1 at the end of the calculation. However it is not guaranteed that the series E(λ) has a radius of convergence |λ_{0}| < 1. It means that the series is divergent for the physical system at λ = 1. One can prove that |λ_{0}| can be obtained by extending λ in the complex plane and looking for the singularities of E(λ). The radius of convergence of the expansion of a function is equal to the distance of the closest singularities in the complex plane to the origin. These singularities are nothing but EPs at λ_{0} and λ_{0}^{∗}. Indeed, for a Hermitian Hamiltonian the singularities occur in complex conjugate pairs with non-zero imaginary parts. The discovery of a partitioning of the Hamiltonian that allowed chemists to get access to a part of the correlation energy using perturbation theory has been a major step in the development of post-Hartree-Fock methods. This case of the Rayleigh-Schrödinger perturbation theory is called the Møller-Plesset perturbation theory [28]. In the MPPT the unperturbed Hamiltonian is the sum of the n monoelectronic Fock operators which are the sum of the one-electron core Hamiltonian h(i), the Coulomb J_{j}(i) and Exchange K_{j}(i) operators.

H_{0} = ∑_{i=1}^{n} f(i)  (4)

f(i) = h(i) + ∑_{j=1,j≠i}^{n} [J_{j}(i) − K_{j}(i)]  (5)

In Hartree-Fock theory you approximate the exact wave function as a Slater-determinant (which is an anti-symmetric combination of mono-electronic orbitals) and those wave functions are eigenvectors of the Fock operators. In the perturbation theory the energy is a power series of λ and the physical energy is obtained by taking λ equal to 1. We will refer to there energy up to the n-th order as the MPn energy. The MP1 energy is the Hartree-Fock energy and the correlation energy is getting recovered from the MP2 level.

E_{MP_{n}} = ∑_{k=0}^{n} E^{(k)}  (6)

But a priori there is no reason that this power series is always convergent for λ=1 when n goes to infinity. In fact, it is known that when the Hartree-Fock wave function is a bad approximation of the exact wave function, for example for multi-reference states, the Møller-Plesset will give bad results[29, 30, 31, 32]. A smart way to investigate the convergence properties of the MP series is to transform the coupling parameter λ into a complex variable. By doing this the Hamiltonian and the energy become functions of this variable. So by searching the singularities of the function E(λ) we can get information on the convergence properties of the MPPT. Those singularities of the energy are exactly the exceptional points connecting the electronic states mentioned in the introduction. The direct computation of the terms of the series is relatively easy up to the 4th order and the 5th and 6th order can be obtained at high cost. But to understand deeply the behavior of the MP series and how it is connected to the singularities, we need to have access to high order terms of the series. For small systems we can have access to the whole series using Full Configuration Interaction. If you diagonalize the Hamiltonian H(λ) in the FCI basis you get the exact energies (in this finite basis set) and expanding in λ allow you to get the Møller-Plesset perturbation series at every order.

3 Historical overview

3.1 Behavior of the Møller-Plesset series

When you use Møller-Plesset perturbation theory it would be very convenient that each time you compute a term of higher order the result obtained is closer to exact energy. In other words that the Mø]ller-Plesset series would be monotonically convergent. Assuming this the only limiting process to get the exact correlation energy in a finite basis is our ability to compute the terms of the perturbation series. Unfortunately this not true in generic cases and rapidly some strange behaviors of the series have been exhibited. In the late 80’s Gill et al. have reported deceptive and slow convergences in stretch systems[29, 30, 31, 32]. In the figure below we can see that the restricted Møller-Plesset series is convergent but oscillating which is not convenient if you are able to compute only few terms (for example here RMP5 is worse than RMP4). On the other hand, the unrestricted Møller-Plesset series is monotonically converging (except for the first few orders) but very slowly so we can’t use it for systems where we can compute only the first terms.

Figure 1: Barriers to homolytic fission of He_{2}^{2+} at MPn/STO-3G level (n = 1–20)[29].

When a bond is stretched the exact function can undergo a symmetry breaking becoming multireference during this process (see for example the case of H_{2} in [33]). A restricted HF Slater determinant is a poor approximation of a broken symmetry wave function but even in the unrestricted formalism, where the spatial orbitals of electrons α and β are not restricted to be the same[34], which allow a better description of symmetry broken system the series doesn’t give accurate results at low orders. Even with this improvement of the zeroth order wave function the series doesn’t have the smooth and rapidly converging behavior wanted. In the unrestricted framework the ground state singlet wave function is allowed to mix with triplet states which lead to spin contamination. Gill et al. highlighted that there is a link between the slow convergence of the unrestricted MP series and the spin contamination of the wave function as it’s shown below within the example of H_{2} in a minimal basis. Handy and co-workers. exhibited the same behavior of the series (oscilating and monotonically slowly) in stretched H_{2}O and NH_{2} systems [31]. Cremer and He performed the same analysis with 29 FCI systems [35] and regrouped all the systems in two classes. The class A systems which have a monotonic convergence to the FCI value and the class B which converge erraticly after initial oscillations. The sample of systems contains stretched molecules and also some at equilibrium geometry, there are also some systems in various basis sets. They highlighted that systems with class A convergence had well-separated electrons pairs whereas class B systems present electrons clustering.

r UHF UMP2 UMP3 UMP4 ⟨S^{2}⟩
0.75 0.0% 63.8% 87.4% 95.9% 0.00
1.35 0.0% 15.2% 26.1% 34.9% 0.49
2.00 0.0% 01.0% 01.8% 02.6% 0.95
2.50 0.0% 00.1% 00.3% 00.4% 0.99

Table 1: Percentage of electron correlation energy recovered and ⟨S^{2}⟩ for the H_{2} molecule as a function of bond length (r,Å) in the STO-3G basis set [30].

This problem was preoccupying in order to develop a method that would allow to compute exact energies for all systems using perturbation theory. But in fact it could be expected as it was known that the perturbation theory was not working well if the wave functions eigenvectors of H_{0} are bad approximations of the exact wave functions. But Olsen et al. have discovered even more preoccupying behavior of the MP series in the late 90’s. They have shown that the series could be divergent even in systems that they considered as well understood like Ne and HF.

Figure 2: Correlation contributions for Ne and HF in the cc-pVTZ-(f/d) ∘ and aug-cc-pVDZ • basis sets.

When the basis sets is augmented with diffuse functions the series are becoming divergent. This is a really worrying observation because chemists want to work in the largest basis set possible to model the complete basis set. They have shown that those divergences were due to a backdoor intruder state i.e. a pair of singularity in the negative half plane. Their method to investigate links between singularities and those behaviors of the series is to scan energies on the real axis to list the avoided crossings. An avoided crossing on the real axis is indicative of a pair of complex conjugate singularities, their real parts are equal to the value of λ at the minimum difference of energy of the avoided crossing. Then modeling an avoided crossing with a two level model, they get an approximation of the pair of singularities by finding the EPs of the two-state model. When the basis set is augmented with diffuse functions the ground state undergoes a shark avoided crossing with a diffuse state on the negative real axis. When there are enough diffuse function in the basis set this pair of singularities is in the unit circle causing the divergence of the series. Olsen and his coworkers conclude that the divergence of the series in those case was due to the interaction of the ground state with a highly diffuse state. They also analyzed the convergence of the series for the CH_{2} as an example molecule containing low-lying doubly excited states. In this molecule the doubly excited states are strongly interacting with the ground state because they have the same spatial and spin symmetry. Thus the single reference wave function is a poor approximation. They have shown that there is a singularity close to the unit circle resulting from an interaction with the lowest doubly excited states and that this singularity is causing the slow convergence for the series of CH_{2}.

3.2 The singularity structure

In the 2000’s Sergeev and Goodson analyzed this problem from a more mathematical point of view by looking at the whole singularity structure where Olsen and his co-workers were trying to find the singularity causing the divergence. They regroup singularities in two classes: the α singularities which have unit order imaginary parts and the β singularities which have very small imaginary parts. The singularities α are related to large avoided crossing between the ground state and a low-lying excited states. Whereas the singularities β come from a sharp avoided crossing between the ground state and a highly diffuse state. They used the work of Stillinger to explain the physical signification of those β singularities. The Møller-Plesset Hamiltonian is defined as above and by reassembling the term we get the expression (10).

H(λ) = H_{0} + λ(H_{phys} − H_{0})  (7)

H_{phys} = ∑_{j=1}^{n} [−\frac{1}{2}∇_{j}^{2} − ∑_{k=1}^{N} \frac{Z_{k}}{|r_{j} − R_{k}|} + ∑_{j<l}^{n} \frac{1}{|r_{j} − r_{l}|}]  (8)

H_{0} = ∑_{j=1}^{n} [−\frac{1}{2}∇_{j}^{2} − ∑_{k=1}^{N} \frac{Z_{k}}{|r_{j} − R_{k}|} + V_{j}^{(scf)}]  (9)

H(λ) = ∑_{j=1}^{n} [−\frac{1}{2}∇_{j}^{2} − ∑_{k=1}^{N} \frac{Z_{k}}{|r_{j} − R_{k}|} + (1 − λ)V_{j}^{(scf)} + λ ∑_{j<l}^{n} \frac{1}{|r_{j} − r_{l}|}]  (10)

The first two terms, the kinetic energy and the electron-nucleus attraction, form the mono-electronic core Hamiltonian which is independant of λ. The third term is the mean field repulsion of the Hartree-Fock calculation done to get H_{0} and the last term is the Coulomb repulsion. If λ is negative, the Coulomb interaction becomes attractive but the mean field stay repulsive as it is proportional to (1 − λ). If λ becomes more and more negative the mean field becomes more and more repulsive so the nucleus can’t bind anymore the electrons because the electron-nucleus attraction is not scaled with λ. But the repulsive mean field is localized around nucleus whereas the electrons interactions persist away from nucleus. So there is a real negative value λ_{c} where the electron form a bound cluster and goes to infinity. According to Baker this value is a critical point of the system and by analogy with thermodynamics the energy E(λ) exhibits a singularity at λ_{c}. Beyond λ_{c} there is a continuum of eigenstates with electrons dissociated from the nucleus. This reasoning is done on the exact Hamiltonian and energy, this is the exact energy which exhibits this singularity on the negative real axis. But in finite basis set, one can prove that for a Hermitian Hamiltonian the singularities of E(λ) occurs in complex conjugate pair with non-zero imaginary parts. Sergeev and Goodson proved that, as predicted by Stillinger, in a finite basis set the critical point on the real axis modeled by a cluster of sharp avoided crossings with diffuse functions, equivalently by a cluster of β singularities in the negative half plane. Finally, it has been shown that β singularities are very sensitive to the basis sets but not to stretching of the system. On the contrary α singularities are relatively insensitive to the basis sets but very sensitive to bond stretching. According to Goodson the singularity structure from molecules stretched from the equilibrium geometry is difficult. To our knowledge the effect of bond stretching on singularities, its link with spin contamination and symmetry breaking of the wave function haven’t been as well understood as the ionization effect and its link with diffuse function. In this work we try to improve our understanding of the effect of symmetry breaking on the singularities of E(λ) and we hope that it will lead to a deeper understanding of perturbation theory.

Perturbation theory in the complex plane  Antoine MARIE

3.3 The physics of quantum phase transition

4 The spherium model References

[1] S. Bittner, B. Dietz, U. Günther, H. L. Harney, M. Miski-Oglu, A. Richter, and F. Schäfer. PT Symmetry and Spontaneous Symmetry Breaking in a Microwave Billiard. Phys. Rev. Lett., 108(2):024101, January 2012.

[2] Y. D. Chong, Li Ge, and A. Douglas Stone. Pt symmetry breaking and laser absorber modes in optical scattering systems. Phys. Rev. Lett., 106(9):093902, March 2011.

[3] N. M. Chtchelkatchev, A. A. Golubov, T. I. Baturina, and V. M. Vinokur. Stimulation of the fluctuation superconductivity by pt symmetry. Phys. Rev. Lett., 109(15):150405, October 2012.

[4] Jörg Doppler, Alexei A. Mailybaev, Julian Böhm, Ulrich Kuhl, Adrian Girschik, Florian Libisch, Thomas J. Milburn, Peter Rabl, Nimrod Moiseyev, and Stefan Rotter. Dynamically encircling an exceptional point for asymmetric mode switching. Nature, 537(7618):76–79, September 2016.

[5] A. Guo, G. J. Salamo, D. Duchesne, R. Morandotti, M. Volatier-Ravat, V. Aimez, G. A. Siviloglou, and D. N. Christodoulides. Observation of pt symmetry breaking in complex optical potentials. Phys. Rev. Lett., 103(9):093902, August 2009.

[6] Chao Hang, Guoxiang Huang, and Vladimir V. Konotop. Pt symmetry with a system of three-level atoms. Phys. Rev. Lett., 110(8):083604, February 2013.

[7] M. Liertzer, Li Ge, A. Cerjan, A. D. Stone, H. E. Türeci, and S. Rotter. Pump-induced exceptional points in lasers. Phys. Rev. Lett., 108(17):173901, April 2012.

[8] Stefano Longhi. Optical realization of relativistic non-hermitian quantum mechanics. Phys. Rev. Lett., 105(1):013903, June 2010.

[9] B. Peng, . K. Ozdemir, S. Rotter, H. Yilmaz, M. Liertzer, F. Monifi, C. M. Bender, F. Nori, and L. Yang. Loss-induced suppression and revival of lasing. Science, 346(6207):328–332, October 2014.

[10] Bo Peng, Şahin Kaya Özdemir, Fuchuan Lei, Faraz Monifi, Mariagiovanna Gianfreda, Gui Lu Long, Shanhui Fan, Franco Nori, Carl M. Bender, and Lan Yang. Parity–time-symmetric whispering-gallery microcavities. Nat. Phys., 10(5):394–398, May 2014.

[11] Alois Regensburger, Christoph Bersch, Mohammad-Ali Miri, Georgy Onishchukov, Demetrios N. Christodoulides, and Ulf Peschel. Parity–time synthetic photonic lattices. Nature, 488(7410):167– 171, August 2012.

[12] Christian E. Rüter, Konstantinos G. Makris, Ramy El-Ganainy, Demetrios N. Christodoulides, Mordechai Segev, and Detlef Kip. Observation of parity–time symmetry in optics. Nat. Phys., 6(3):192–195, March 2010.

[13] Joseph Schindler, Ang Li, Mei C. Zheng, F. M. Ellis, and Tsampikos Kottos. Experimental study of active LRC circuits with pt symmetries. Phys. Rev. A, 84(4):040101, October 2011.

[14] Alexander Szameit, Mikael C. Rechtsman, Omri Bahat-Treidel, and Mordechai Segev. Ptsymmetry in honeycomb photonic lattices. Phys. Rev. A, 84(2):021806, August 2011.

[15] K. F. Zhao, M. Schaden, and Z. Wu. Enhanced magnetic resonance signal of spin-polarized Rb atoms near surfaces of coated cells. Phys. Rev. A, 81(4):042903, April 2010.

[16] C. Zheng, L. Hao, and G. L. Long. Observation of a fast evolution in a parity-time-symmetric system. Philos. Trans. R. Soc. Math. Phys. Eng. Sci., 371(1989):20120053–20120053, March 2013.

[17] Youngsun Choi, Choloong Hahn, Jae Woong Yoon, and Seok Ho Song. Observation of an antipt-symmetric exceptional point and energy-difference conserving dynamics in electrical circuit resonators. Nat. Commun., 9(1):2182, December 2018.

[18] Ramy El-Ganainy, Konstantinos G. Makris, Mercedeh Khajavikhan, Ziad H. Musslimani, Stefan Rotter, and Demetrios N. Christodoulides. Non-hermitian physics and pt symmetry. Nat. Phys., 14(1):11–19, January 2018.

[19] W D Heiss and A L Sannino. Avoided level crossing and exceptional points. J. Phys. Math. Gen., 23(7):1167–1178, April 1990.

[20] W. D. Heiss. Phases of wave functions and level repulsion. Eur. Phys. J. D, 7:1, 1999.

[21] W D Heiss. The physics of exceptional points. J. Phys. Math. Theor., 45(44):444016, November 2012.

[22] Dieter Heiss. Circling exceptional points. Nat. Phys., 12:823–824, 2016.

[23] D. R. Yarkony. Diabolical conical intersections. Rev. Mod. Phys., 68:985, 1996.

[24] M. V. Berry. Quantal Phase Factors Accompanying Adiabatic Changes. Proc. Royal Soc. A, 392:45, 1984.

[25] N. Moiseyev. Non-Hermitian Quantum Mechanics. Cambridge University Press, 2011.

[26] Jeppe Olsen, Ove Christiansen, Henrik Koch, and Poul Jørgensen. Surprising cases of divergent behavior in møller-plesset perturbation theory. J. Chem. Phys., 105:5082–5090.

[27] Jeppe Olsen, Poul Jørgensen, Trygve Helgaker, and Ove Christiansen. Divergence in møller–plesset theory: A simple explanation based on a two-state model. J. Chem. Phys., 112(22):9736–9748.

[28] Chr. Møller and M. S. Plesset. Note on an approximation treatment for many-electron systems. Phys. Rev., 46(7):618–622.

[29] Peter M. W. Gill and Leo Radom. Deceptive convergence in møller-plesset perturbation energies. Chemical Physics Letters, 132(1):16–22.

[30] Peter M. W. Gill, John A. Pople, Leo Radom, and Ross H. Nobes. Why does unrestricted mo/ller–plesset perturbation theory converge so slowly for spin-contaminated wave functions? J. Chem. Phys., 89(12):7307–7314.

[31] N. C. Handy, P. J. Knowles, and K. Somasundram. On the convergence of the møller-plesset perturbation series. Theoret. Chim. Acta, 68(1):87–100.

[32] M. B. Lepetit, M. Pélissier, and J. P. Malrieu. Origins of the poor convergence of many-body perturbation theory expansions from unrestricted hartree–fock zeroth-order descriptions. J. Chem. Phys., 89(2):998–1008.

[33] A. Szabo and N. S. Ostlund. Modern quantum chemistry. McGraw-Hill, New York, 1989.

[34] Hideo Fukutome. Unrestricted hartree–fock theory and its applications to molecules and chemical reactions. 20(5):955–1065.

[35] D. Cremer and Z. He. Sixth-order møllerplesset perturbation theory on the convergence of the mpn series. J. Phys. Chem., 100(15):6173––6188.